### v2

# MissionGNN: Hierarchical Multimodal GNN-based Weakly Supervised Video Anomaly Recognition with Mission-Specific Knowledge Graph Generation

Abstract
In the context of escalating safety concerns across various domains, the tasks of Video Anomaly Detection (VAD) and Video Anomaly Recognition (VAR) have emerged as critically important for applications in intelligent surveillance, evidence investigation, violence alerting, etc. These tasks, aimed at identifying and classifying deviations from normal behavior in video data, face significant challenges due to the rarity of anomalies which leads to extremely imbalanced data and the impracticality of extensive frame-level data annotation for supervised learning. This paper introduces a novel hierarchical graph neural network (GNN) based model MissionGNN that addresses these challenges by leveraging a state-of-the-art large language model and a comprehensive knowledge graph for efficient weakly supervised learning in VAR. Our approach circumvents the limitations of previous methods by avoiding heavy gradient computations on large multimodal models and enabling fully frame-level training without fixed video segmentation. Utilizing automated, mission-specific knowledge graph generation, our model provides a practical and efficient solution for real-time video analysis without the constraints of previous segmentation-based or multimodal approaches. Experimental validation on benchmark datasets demonstrates our model’s performance in VAD and VAR, highlighting its potential to redefine the landscape of anomaly detection and recognition in video surveillance systems.
I Introduction
The escalating concern for safety across various sectors has heightened the importance of tasks like Video Anomaly Detection (VAD) and Video Anomaly Recognition (VAR). These tasks are crucial for applications such as intelligent surveillance [3, 4], evidence investigation [19], and violence alerting [23, 10], where the goal is to automatically identify and classify activities that deviate from normal patterns in video data [26]. VAD and VAR can potentially be approached through supervised learning to detect anomalies [2, 33]; however, the rarity of such anomalies leads to highly imbalanced datasets, complicating the training process. Moreover, the extensive effort and cost required for frame-level video data annotation render supervised learning with frame-level annotations impractical. Consequently, many researchers have shifted their focus towards weakly-supervised learning [13, 27, 29, 37, 42, 49], which relies on video-level annotations to train models. This shift addresses the challenges posed by data imbalance and annotation costs but introduces new complexities in achieving accurate frame-level anomaly detection and recognition.
Recent advancements in tackling these challenges involve the use of pre-trained models to leverage the rich, embedded information they offer, thereby enhancing the distinction between normal and abnormal frames [42, 49, 27, 35, 21]. However, for VAR, the intricacies go beyond what pre-trained vision models can discern, necessitating finer granularity. To address this, some studies [42, 35, 34, 39, 21] have explored large multimodal neural networks [22, 24] capable of extracting more nuanced distributions, employing methods like learnable prompts [48] or adapters [35] to refine model differentiation. This approach, while innovative, demands substantial computational resources due to the gradient flow through these large models [9]. Furthermore, it is incompatible with the prevailing trend towards large pre-trained models, which are typically accessed via APIs, making it impossible to compute gradients directly.
Moreover, the application of multiple-instance learning (MIL) with fixed video segmentation has been a popular simplified strategy among researchers [27]. By dividing videos into fixed-size or fixed-number segments and introducing a transformer model to capture not only short-term relationships but also long-term relationships between segments, these models have shown improved performance [42, 35, 46, 29, 13, 37]. Yet, this method struggles with the variability of anomaly event durations e.g., events that happen in very short periods or video length is much longer than the event length, and the practical limitations of real-time analysis, where future frame data is unavailable, and past frame data is potentially infinite. Such reliance on long-term relationship modeling further complicates the deployment of these models in real-world scenarios, where rapid and continuous analysis is essential.
This paper proposes a novel approach MissionGNN to tackle VAR by introducing a novel hierarchical graph neural network (GNN) based model that avoids the computational burdens associated with gradient computation of large multimodal models and introduces a fully frame-level training methodology without using MIL. Unlike prior work that limitedly utilized GNNs for temporal information processing [46, 30], our model uses GNN to capture semantic information using a mission-specific knowledge graph (KG) that is automatically generated by our KG generation framework using a large language model namely GPT-4 [1], and a large knowledge graph namely ConceptNet [25]. Instead of flowing gradient to a large multimodal model, our approach utilizes a built mission-specific knowledge graph and only uses the large multimodal model without any gradient computation to have initial embeddings for each node that will be used in our GNN. After that, our approach uses a small transformer to capture only short-term relationships. This innovative strategy enables effective anomaly detection without the need for gradient computations on large models, relying instead on light models while offering a practical, efficient, and scalable solution to the challenges of weakly-supervised learning in video anomaly recognition.
In summary, our work represents a fundamentally novel contribution to the field, offering the following key advancements:
  • To the best of our knowledge, we propose MissionGNN a novel hierarchical GNN-based model with an automated mission-specific knowledge graph generation framework to have an efficient weakly supervised video anomaly recognition model. Through our proposed mission-specific knowledge graph generation framework, our approach can be adopted in various other tasks.
  • Unlike previous works on using pre-trained large multimodal models, our approach does not require heavy gradient computation on the pre-trained large multimodal models providing efficient training. At the same time, our approach provides additional flexibility that also can be implemented using large pre-trained models that are exclusively accessible through APIs which we do not have access to the weights of the model.
  • Inspired by human intuition, we introduce completely frame-level training and inference that does not require video segmentation which can affect the model’s performance by merging multiple frames into fixed-size segments. Furthermore, this makes our model more practical in real applications by not using future segments or past segments suitable for real-time applications.
II Related Works
II-A Weakly Supervised Video Anomaly Detection and Recognition
Video anomaly detection has traditionally relied on fully supervised methods requiring frame-level annotations, which are costly and time-consuming [2, 33]. To address these challenges, weakly supervised approaches using video-level labels [13, 27, 29, 37], one-class classification [16, 18], and unsupervised methods [20, 41] have gained traction. Weakly supervised techniques, in particular, offer a balance between performance and annotation costs, making them popular in recent research.
MIL [27], which divides videos into segments and labels them uniformly, is a key method in weakly supervised video anomaly detection. Enhancements like internal loss constraints [44] and pseudo labeling [46] have enabled models to capture both short- and long-term relationships [29, 13]. However, MIL struggles with variable anomaly durations and is impractical for real-time scenarios due to its reliance on long-term dependencies, which our work seeks to overcome its limitations.
While VAD has been widely studied, VAR is less explored due to the scarcity of labeled anomalies [27]. Recent efforts to apply Large Language and Vision (LLV) models [22, 24] in VAR have shown promising results [42, 43]. These models, trained on large image-caption datasets, can improve anomaly recognition. Our work adapts LLV models for weakly supervised VAR, addressing key challenges in this domain.
II-B Graph Neural Networks and Knowledge Graphs
Graph Neural Networks, such as Graph Convolution Networks (GCNs) [45] and Graph Attention Networks (GATs) [32], are widely used to process structural data represented as nodes and edges. The Message Passing Network framework [7] provides a generalized GNN model, where messages are computed for all edges, aggregated by nodes, and updated to generate node embeddings. These embeddings capture the graph’s structural essence and are critical for tasks requiring an understanding of relationships between entities.
GNNs are often combined with knowledge graphs to enhance tasks like question answering [40], information retrieval [5], and reasoning [6]. A key resource in this space is ConceptNet [25], a large-scale semantic knowledge graph that encodes relationships between words, such as “IsA” or “RelatedTo,” across multiple languages. ConceptNet is widely used in various domains to provide structured context and reasoning capabilities. In our work, ConceptNet serves as one of the core components of our automated knowledge graph generation framework, supplying structured knowledge that aids in anomaly recognition.
While previous approaches [46, 30] have used GNNs for VAD, they mainly focused on capturing temporal relationships between video frames. In contrast, our method employs a novel GNN architecture over a knowledge graph constructed through automated generation to perform semantic reasoning for VAR tasks, extending GNN capabilities beyond temporal analysis.
II-C Large Multimodal Models and Prompt Optimization
Figure 1: The framework for mission-specific knowledge graph generation.
Large Language and Vision models [22, 24], which integrate text and image data in a joint embedding space, have proven successful in various tasks, including anomaly detection [39, 35]. These models use modality-specific encoders and contrastive learning to align representations across different modalities. AnomalyCLIP [42], which leverages CLIP [22], demonstrated improved VAR performance by harnessing the strengths of LLV models. Expanding on this, recent models like ImageBind [8] support multiple modalities. Our work builds on this progress by incorporating ImageBind to explore multimodal anomaly detection.
Prompt optimization techniques were introduced to enhance model performance by refining embedding representations. Initial methods relied on manually crafted prompts [11], followed by learnable prompt techniques [14, 12], which automatically optimize prompts using gradients. However, these approaches often face overfitting and scalability issues [47]. To overcome these limitations, our research replaces prompt optimization with knowledge graph reasoning, which acts as knowledge extraction from a large language model, providing a more efficient and scalable approach for improving model efficacy.
III Methodology
III-A Overall Pipeline
Figure 2: The overall framework for our proposed model utilizing the novel concept of hierarchical graph neural network.
Our proposed approach consists of two major frameworks: mission-specific knowledge graph generation and hierarchical graph neural network-based detection. First, the mission-specific KG generation framework is used for generating KG that can extract useful information from a given frame image. In VAR, each mission-specific KG indicates a KG that contains structured knowledge regarding the corresponding anomaly type. Through this automated framework, our method builds $n$ mission-specific KG each corresponding to one of the $n$ anomaly types. The generated mission-specific KGs are used in the following framework, hierarchical GNN-based detection. $n$ hierarchical GNNs are paired with one of the $n$ mission-specific KGs to conduct reasoning over the extracted knowledge to conduct the corresponding mission which is detecting a certain type of anomaly. And large multimodal model’s embedding of each frame is processed by each of the GNN models. Now, embeddings from the final embedding node of each GNN model are concatenated into a single embedding vector that represents reasoning results from all mission-specific KGs. The final vector now is conveyed to a short-term temporal model with the previous consecutive 29 frames. Note that this temporal model only considers short-term relationships which does not require long frame sequence as an input. The resulting output from the short-term temporal model passes through a final decision model to have a final score distribution for each anomaly.
Details are discussed in the following subsections: we first address our mission-specific KG generation framework in subsection III-B, then the proposed hierarchical multimodal GNN model and short-term temporal model are introduced in subsection III-C and subsection III-D respectively, and finally training losses are described in subsection III-E.
III-B Mission-Specific Knowledge Graph Generation Framework
Figure 1 shows the overall framework for mission-specific knowledge graph generation. Given a mission $m_{i}$ from a user, the framework generates a KG $G_{m_{i}}=\left(V_{m_{i}},E_{m_{i}}\right)$, which is directed acyclic graph (DAG). In order to generate $G_{m_{i}}$, it first, combines the given mission $m_{i}$ with a pre-formatted prompt to have a query message. Then, $n_{cpt}$ number of key concept words $w^{cpt}_{m_{i},j}$ are extracted by querying the generated query message to a large language model. Each word $w^{cpt}_{m_{i},j}$ is considered as a key concept node $v^{cpt}_{m_{i},j}$ and those key concept nodes build part of the KG $\{v^{cpt}_{m_{i},j}\}=V^{cpt}_{m_{i}}\subset V_{m_{i}}$. Next, to expand KG, iterative sub-graph extraction is conducted using the large language model. First sub-graph nodes $v^{sub_{1}}_{m_{i},j}\in V^{sub_{1}}_{m_{i}}$ are extracted by querying the key concept nodes to the large language model. A set of edges containing neighboring relationships $E^{sub_{1}}_{m_{i}}=\{\left(v,u\right)\in V^{cpt}_{m_{i}}\times V^{sub_{1}}_{m_{i}}|\left(v,u\right)\in\textbf{ConceptNet}\}$ is also extracted during the process. In our implementation, we only considered the “/r/RelatedTo” relationship type. Next sub-graph nodes $v^{sub_{2}}_{m_{i},j}\in V^{sub_{2}}_{m_{i}}$ are extracted with the similar way but directly neighboring to $V^{sub_{1}}_{m_{i}}$ with edges $E^{sub_{2}}_{m_{i}}=\{\left(v,u\right)\in V^{sub_{1}}_{m_{i}}\times V^{sub_{2}}_{m_{i}}|\left(v,u\right)\in\textbf{ConceptNet}\}$. This process, which is extracting $k+1^{th}$ set of sub-graph nodes $V^{sub_{k+1}}_{m_{i}}$ and edges $E^{sub_{k+1}}_{m_{i}}$ from $V^{sub_{k}}_{m_{i}}$, is iteratively conducted until $V^{sub_{d_{sub}}}_{m_{i}}$ and $E^{sub_{d_{sub}}}_{m_{i}}$ are retrieved where $d_{sub}$ indicates maximum depth of the sub-graph. The extracted sub-graph nodes $V^{sub}_{m_{i}}=\bigcup_{1\leq k\leq d_{sub}}{V^{sub_{k}}_{m_{i}}}$ and edges $E^{sub}_{m_{i}}=\bigcup_{1\leq k\leq d_{sub}}{E^{sub_{k}}_{m_{i}}}$ also build part of the KG by $V^{sub}_{m_{i}}\subset V_{m_{i}}$ and $E^{sub}_{m_{i}}\subset E_{m_{i}}$. Finally, $G_{m_{i}}$ is completed by combining sensor node $v^{snr}_{m_{i}}$ and encoding node $v^{ecd}_{m_{i}}$. $v^{snr}_{m_{i}}$ is attached to the graph by connecting it with key concept nodes $v^{cpt}_{m_{i},j}$ towards $v^{cpt}_{m_{i},j}$ direction while $v^{ecd}_{m_{i}}$ is attached by connecting it with leaf nodes, which are nodes that do not have outgoing edges, of sub-graph nodes towards $v^{ecd}_{m_{i}}$ direction. Completed $G_{m_{i}}$ is described in Equation 1.
$\begin{split}V_{m_{i}}=&\left\{v^{snr}_{m_{i}}\right\}\cup V^{cpt}_{m_{i}}\cup V^{sub}_{m_{i}}\cup\left\{v^{ecd}_{m_{i}}\right\}\\ E_{m_{i}}=&\left\{\left(v^{snr}_{m_{i}},u\right)|u\in V^{cpt}_{m_{i}}\right\}\\ &\cup E^{sub}_{m_{i}}\\ &\cup\left\{\left(v,v^{ecd}_{m_{i}}\right)|v\in V^{sub}_{m_{i}}\wedge\left(v,u\right)\notin E^{sub}_{m_{i}}\forall u\in V^{sub}_{m_{i}}\right\}\end{split}$ (1)
To process generated KG $G_{m_{i}}$ using GNN later, it is necessary to have embeddings of the nodes. For having embedding representations, we use text encoder $\mathcal{E}_{T}(.)$ for embedding key concept nodes and sub-graph nodes and image encoder $\mathcal{E}_{I}(.)$ for embedding the sensor node. By using encoders $\mathcal{E}_{T}(.)$ and $\mathcal{E}_{I}(.)$ from the pre-trained ImageBind model, we can have embeddings in the same embedding space for different modalities. For the encoding node, it is initialized with a 00 valued vector. In VAR, multiple mission-specific KGs $G_{m_{i}}$ are generated by setting each mission $m_{i}$ as detecting each anomaly type $T_{i}$. As a result, for the VAR task of $n$ number of anomaly types, a total of $n$ mission-specific KGs $G_{T_{i}}$ are built by the framework.
III-C Hierarchical Multimodal GNN
For a given frame image $F_{t}$ and mission-specific KGs $G_{T_{i}}$, to conduct reasoning for each KG, our proposed hierarchical multimodal GNN model is used over each given KG as presented in Figure 2. For each KG $G_{T_{i}}$, the reasoning process propagates from a sensor node of $G_{T_{i}}$ to its embedding node. To conduct the reasoning process, first, a given frame image $F_{t}$ is encoded by $\mathcal{E}_{I}(F_{t})$ and used as an embedding vector of the sensor node $v^{snr}_{m_{i}}$. Next, $d_{sub}+2$ GNN layers are applied hierarchical way so that reasoning by GNN layers can propagate starting from the sensor node and the final resulting embedding can be located in the embedding node. Each GNN layer $\mathcal{G}_{m_{i},l}(.)$ consists of 5 sub-layers: dense layer, hierarchical message passing layer, hierarchical aggregate layer, batch normalization layer, and activation layer. Dense layer $\phi_{m_{i},l}\colon\mathbb{R}^{|V_{m_{i}}|\times D_{l-1}}\to\mathbb{R}^{|V_{m_{i}}|\times D_{l}}$, where $D_{m_{i},l-1}$ indicates dimensionality of previous layer and $D_{m_{i},l}$ indicates dimensionality of the current layer, transfers embedding space to have better relationship representations between each knowledge embedding and can be formulated as follows:
$\begin{split}\phi_{m_{i},l}(X_{l-1})=\mathbf{W}^{\phi}_{m_{i},l}X_{l-1}+\mathbf{b}^{\phi}_{m_{i},l}\end{split}$ (2)
where $\mathbf{W}^{\phi}_{m_{i},l}$ is trainable parametric matrix, $\mathbf{b}^{\phi}_{m_{i},l}$ is trainable bias, and $X_{l-1}\in\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l-1}}$ is embeddings of previous layer for nodes. Hierarchical message passing layer $\mathcal{M}^{h}_{m_{i},l}\colon\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}}\to\mathbb{R}^{|E^{(l)}_{m_{i}}|\times D_{m_{i},l}}$ compute messages each corresponding to each node where $E^{(l)}_{m_{i}}\subset E_{m_{i}}$ indicates a set of edges that connect depth $l-1$ nodes $V^{(l-1)}_{m_{i}}\subset V_{m_{i}}$ and depth $l$ nodes $V^{(l)}_{m_{i}}\subset V_{m_{i}}$. Note that depth 00 node set $V^{(0)}_{m_{i}}$ contains only one node which is sensor node $v^{snr}_{m_{i}}$ and $l$ starts from $1$. Each message indicates measured relativity by the element-wise product between two knowledge embeddings as described below:
$\begin{split}\mathcal{M}^{h}_{m_{i},l}(X)=\{X_{s}\cdot X_{d}\}_{(s,d)\in E^{(l)}_{m_{i}}}\end{split}$ (3)
where $s$ indicates source node, $d$ indicates destination node, and $X\in\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}}$ is node embeddings. Hierarchical aggregate layer $\mathcal{A}^{h}_{m_{i},l}\colon\{\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}},\mathbb{R}^{|E^{(l)}_{m_{i}}|\times D_{m_{i},l}}\}\to\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}}$ aggregates all the passed messages from source nodes into a single vector by getting a mean vector meanwhile remaining other embeddings of nodes that do not have messages from their source nodes, denoted as:
$\begin{split}&\mathcal{A}^{h}_{m_{i},l}(X,M)=\\ &\left\{X_{d}\mathbf{1}(d\notin V^{(l)}_{m_{i}})+\frac{\sum_{(s,d)\in E^{(l)}_{m_{i}}}{M_{s,d}}}{|\{s\in V_{m_{i}}|(s,d)\in E^{(l)}_{m_{i}}\}|}\mathbf{1}(d\in V^{(l)}_{m_{i}})\right\}_{d\in V_{m_{i}}}\end{split}$ (4)
where $X\in\mathbb{R}^{|V_{m_{i}}|\times D_{m_{i},l}}$ indicates node embeddings, $M\in\mathbb{R}^{|E^{(l)}_{m_{i}}|\times D_{m_{i},l}}$ indicates messages from $\mathcal{A}^{h}_{m_{i},l}$, and $\mathbf{1}(.)$ indicates a function that gives $1$ if a given statement is true otherwise 00. The final form of the GNN layer at layer $l$ is formulated as follows:
$\begin{split}&\mathcal{G}_{m_{i},l}(X_{l-1})=\text{ELU}(\\ &\text{BatchNorm}\left(\mathcal{A}^{h}_{m_{i},l}\left(\phi_{m_{i},l}\left(X_{l-1}\right),{M}^{h}_{m_{i},l}\left(\phi_{m_{i},l}\left(X_{l-1}\right)\right)\right)\right))\end{split}$ (5)
[…]
III-D Short-term Temporal Model
Using multiple KGs and reasoning over them with our proposed hierarchical GNNs only considers information from a single frame. To extract reacher information by exploiting video data, we use transformer-based temporal model $\mathcal{T}\colon\mathbb{R}^{T\times D}\to\mathbb{R}^{D}$ where $T$ indicates the number of consecutive frames. Specifically, we use simple a single transformer encoder layer [31] and put a sequence of reasoning embeddings including reasoning embeddings of previous $T-1$ consecutive frames $\vec{F}_{t}=\{\vec{f}_{t-T+1},\vec{f}_{t-T+2},\cdots,\vec{f}_{t}\}\in\mathbb{R}^{T\times D}$ to the transformer and only takes the last output embedding $\vec{f}^{\prime}_{t}=\mathcal{T}(\vec{F}_{t})\in\mathbb{R}^{D}$ which is corresponding to the last input $\vec{f}_{t}$. This process is also presented in Figure 2. Unlike previous methods that considered not only short-term temporal relationships but also long-term temporal relationships, our model focuses on short-term relationships only by setting $T=30$ without any video segmentation. In VAR, we assume that anomalies occur in short time intervals by following human intuition, i.e. when people encounter anomalies such as gunshots, they can detect them immediately without having to understand long-term relationships. The output of the short-term temporal model $\vec{f}^{\prime}_{t}$ is then processed by a decision model $f^{dec}\colon\mathbb{R}^{D}\to\mathbb{R}^{n+1}$ which is consists of a simple linear layer and softmax function:
$\begin{split}f^{dec}(\vec{f}^{\prime}_{t})=\text{Softmax}(\mathbf{W}^{dec}\vec{f}^{\prime}_{t}+\mathbf{b}^{dec})\end{split}$ (6)
where $\mathbf{W}^{dec}$ is trainable parametric matrix, and $\mathbf{b}^{dec}$ is trainable bias. The output of the decision model $\vec{s}_{t}=f^{dec}(\vec{f}^{\prime}_{t})$ contains the estimated probability of the given frame $F_{t}$ is normal $p_{N}(F_{t})=\vec{s}_{t,0}$ and probabilities of the frame belong to each anomaly type $i\in\{1,2,\cdots,n\}$: $p_{A,i}(F_{t})=p_{A}(F_{t})p_{i|A}(F_{t})=\vec{s}_{t,i}$ where $p_{A}(F_{t})=1-p_{N}(F_{t})$ indicates probability of abnormal and $p_{i|A}(F_{t})=\frac{\vec{s}_{t,i}}{1-p_{N}(F_{t})}$ indicates conditional probability of anomaly type $i$ when it is anomaly.
III-E Weakly-supervised Frame-level Training
Decaying Threshold-based Anomaly Localization. A better performing model requires estimating normal frames and anomaly frames among frames in the same video, where a single type of anomaly is labeled in weakly-supervised learning. However, it is difficult to implement a widely used rank-based approach without using MIL-based training. To resolve this issue and have a model that is capable of fully frame-level training and inference, we introduce a novel anomaly localization method, which is a method that separates normal frames from anomaly-labeled video, using our proposed concept of decaying threshold. Decaying threshold $\theta_{d}$ is denoted as follows:
$\begin{split}\theta^{(iter)}_{d}=\frac{\theta_{0}\alpha^{iter}_{d}}{2}+0.5\end{split}$ (7)
where $\theta_{0}$ indicates initial value of $\theta_{d}$, $\alpha_{d}$ indicates threshold decay rate, and $iter$ indicates current iteration of training. For each training iteration $iter$, frames $F_{t}$ from anomaly-labeled videos are considered as normal frames when $p_{A}(F_{t})=1-p_{N}(F_{t})\leq 1-\theta^{(iter)}_{d}$ is satisfied. In our implementation, we set $\theta_{0}=1$ so that the model initially considers every frame in anomaly-labeled videos as an anomaly frame and slowly approaches $0.5$ as it separates normal frames. We assume that while all reasoning embeddings of frames $F^{A}_{t}$ from anomaly-labeled videos separated from reasoning embeddings of frames $F^{N}_{t}$ from normal-labeled videos, the ones $F^{A+}_{t}$ that are similar to normal-labeled frames ($p_{A}(F^{A+}_{t})\leq 1-\theta^{(iter)}_{d}$) slowly separated from the anomaly distribution, which makes them be separated later from anomaly distribution $F^{A-}_{t}$ that separated faster ($p_{A}(F^{A-}_{t})>1-\theta^{(iter)}_{d}$) by the decaying threshold.
Cross-entropy Loss. After the decaying threshold-based anomaly localization, we have normal frames $F^{N}_{t}$, normal estimated frames $F^{A+}_{t}$, and anomaly estimated frames $F^{A-}_{t}$. We use the cross-entropy loss to maximize probabilities that match each case. Thus, we have three cross-entropy terms for each set of cases as follows:
$\mathcal{L}_{N}$ $=-\frac{\lambda_{N}}{|F^{N}|}\sum_{t}{\log{p_{N}(F^{N}_{t})}}$ (8)
$\mathcal{L}_{A+}$ $=-\frac{\lambda_{N}}{|F^{A+}|}\sum_{t}{\log{p_{N}(F^{A+}_{t})}}$ (9)
$\mathcal{L}_{A-}$ $=-\frac{1}{|F^{A-}|}\sum_{t}{\lambda_{A,y_{t}}\log{p_{A,y_{t}}(F^{A-}_{t})}}$ (10)
where $y_{t}$ indicates the corresponding anomaly type, $\lambda_{N}$ indicates loss weight for normal cross-entropy, and $\lambda_{A,c}$ indicates loss weight for anomaly class $c$ cross-entropy. $\lambda_{N}$ and $\lambda_{A,c}$ are adjusted to cope with data imbalance.
Sparsity and Smoothing Losses. We adapted the same two regularization loss terms that are used in [28]. First is done through sparsity loss term exploiting that the anomaly cases are rare in entire time frames. Sparsity loss tries to minimize the number of frames that are predicted as anomalies by following:
$\mathcal{L}_{spa}=\frac{1}{|F^{A}|}\sum_{t}{p_{A}(F^{A}_{t})}$ (11)
Note that it is only applied to frames from anomaly-labeled videos. Also, to make predictions smoother along the time dimension, another loss term smoothing loss is implemented as follows:
$\mathcal{L}_{smt}=\frac{1}{|F|}\sum_{t}{\left(p_{A}(F_{t})-p_{A}(F_{t-1})\right)^{2}}$ (12)
where $F_{t}$ and $F_{t-1}$ are consecutive two frames.
By combining all loss terms, final loss $\mathcal{L}$ is formed as follows:
$\mathcal{L}=\mathcal{L}_{N}+\mathcal{L}_{A+}+\mathcal{L}_{A-}+\lambda_{spa}\mathcal{L}_{spa}+\lambda_{smt}\mathcal{L}_{smt}$ (13)
where $\lambda_{spa}$ and $\lambda_{smt}$ indicate balance coefficients for sparsity loss and smoothing loss term respectively.
IV Experiments
TABLE I: Memory requirements for computational graph over each model in MiB.
Batch Size Model Efficiency ImageBind MissionGNN 1 2,171 787 2.75×2.75$\times$2.75 × 32 5,240 960 5.45×5.45$\times$5.45 × 64 5,594 1,508 3.70×3.70$\times$3.70 × 128 5,948 2,228 2.66×2.$6$
IV-A Experimental Settings
IV-A1 Implementation Details.
For our multimodal embedding model, we use the pre-trained ImageBind huge model. Also, for the mission-specific KG generation, we use GPT-4 and ConceptNet 5. The loss term balance coefficients $\lambda_{spa}$ and $\lambda_{smt}$ are set to $0.001$ for both of them. We utilize the AdamW [17] as the optimizer with learning rate of $10^{-5}$, a weight decay of $1.0$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. The decaying threshold $\alpha_{d}$ is set to $0.9999$. For our GNN models, we set the dimensionality all the same $D_{m_{i},l}=8$. In the case of the short-term temporal model, we use an inner dimensionality of 128 with 8 attention heads. Finally, the model is trained for 3,000 steps with a mini-batch of 128.
IV-A2 Evaluation Metrics.
We perform evaluations that include both VAD and VAR. Consistent with previous research, VAD performance is measured using the area under the curve (AUC) of the receiver operating characteristics (ROC) at the frame level. A higher AUC at the frame level indicates a better ability to distinguish normal from abnormal events. For the VAR evaluation, we extend the AUC metric to the multi-classification scenario. Specifically, we compute the AUC for each anomalous class, treating anomalous frames of the class as positive instances and all others as negative instances. We then compute the mean AUC (mAUC) across all anomalous classes. Likewise, for the XD-Violence dataset, we follow the established evaluation protocol [38] and VAD results are presented using the average precision (AP) derived from the precision-recall curve (PRC). For VAR outcomes, we report the mean average precision (mAP), which involves averaging the binary AP values across all abnormal classes.
IV-B Datasets
Our experiments utilize two Video Anomaly Detection (VAD) datasets: UCF-Crime [28] and XD-Violence [38]. Additionally, the ShanghaiTech dataset [15], comprising 437 videos captured from diverse surveillance cameras on a university campus, is commonly employed in this research field. However, since some anomaly class labels are only present in the training dataset, we have determined that ShanghaiTech is not suitable for evaluating our method. Therefore, we have chosen not to include it in our evaluation process.
UCF-Crime is an extensive dataset comprising 1900 untrimmed surveillance videos depicting 13 real-world anomalies with significant implications for public safety. The training set comprises 800 normal and 810 anomalous videos, while the testing set includes the remaining 150 normal and 140 anomalous videos.
XD-Violence, a large-scale violence detection dataset, which consists of 4754 untrimmed videos with audio signals and weak labels. It is partitioned into a training set of 3954 videos and a test set of 800 videos, totaling 217 hours in duration and covering various scenarios with 6 anomaly categories. Notably, each violent video may carry multiple labels, ranging from 1 to 3. To align with our training setup, where only one anomaly type per video is considered, we curate a subset comprising 4463 videos containing at most one anomaly.
IV-C Evaluation on Training Memory Efficiency
To demonstrate the advantage in training in terms of memory consumption of our proposed approach, we measured GPU memory consumption specifically the amount of memory for computing computational graphs which is required to conduct gradient backpropagation. Table I shows measured GPU memory consumption used for computational graphs for each of ImageBind and our GNN model by different batch sizes. We can observe that our model consumes 2.66×2.$6$ to 5.45×5.45$\times$5.45 × lower memory. Considering previous approaches used learnable prompts or other types of gradient computation of the LLVs are also using additional models to process the output from the LLV, the memory efficiency can be increased. It enables memory-efficient training while maintaining or outperforming previous models in some benchmarks that are shown in subsection IV-D.
IV-D Evaluation on VAD and VAR
Table II and Table III show the evaluation results of both VAD and VAR tasks on UCF-Crime dataset and XD-Violence dataset respectively. MissionGNN surpasses the performance of cutting-edge VAR tasks in the UCF-Crime dataset, albeit showing a slight lag in VAD tasks. In contrast, AnomalyCLIP exhibits superior performance in VAR tasks within the XD-Violence dataset, while our framework notably outperforms state-of-the-art VAD tasks in the same dataset.
IV-E Evaluation on Real-time VAD and VAR
A key advantage of the MissionGNN framework is its ability to operate with only 30 consecutive frames during both training and inference, making it well-suited for real-time anomaly detection and recognition. In contrast, AnomalyCLIP, which achieves higher performance in some evaluations, requires the entire video sequence for both training and inference, limiting its practical use in real-time applications such as intelligent surveillance. To illustrate this distinction, we conducted a comparative evaluation on the UCF-Crime and XD-Violence datasets, constraining AnomalyCLIP to the same 30-frame input as our model. The results, shown in Table II and Table III, highlight the impact of this limitation on AnomalyCLIP’s performance. In the VAR task, restricting the frame count caused a significant performance drop across all anomaly classes for AnomalyCLIP, resulting in reductions of 4.5% in mAUC and 14.29% in mAP scores across both datasets. This adjustment altered the performance comparison between the two methods. On the UCF-Crime dataset, MissionGNN exhibited a 6.23% improvement over AnomalyCLIP, further widening the performance gap. On the XD-Violence dataset, however, the gap narrowed by 3.71%, reducing the original 18% difference, which emphasizes the limitations of previous approaches in real-time settings. In the VAD task, MissionGNN outperformed AnomalyCLIP on both datasets, with gains of 1.13% in AUC and 32.91% in AP scores. These findings underscore the performance degradation of AnomalyCLIP when adapted for real-time analysis, as reflected in its lower AUC and AP scores on both datasets, reinforcing the practical advantages of our proposed framework for real-time anomaly detection and recognition.
TABLE II: Results of the state-of-the-art methods and MissionGNN on UCF-Crime
Method Class mAUC AUC Abuse Arrest Arson Assault Burglary Explosion Fighting RoadAcc. Robbery Shooting Shoplifting Stealing Vandalism CLIP zero-shot [22] 57.37 80.65 91.72 80.83 74.34 90.31 83.54 87.46 70.22 63.99 71.21 45.49 66.45 74.12 58.63 RTFM [29] 79.99 62.57 90.53 82.27 85.53 92.76 85.21 90.31 81.17 82.82 92.56 90.23 87.20 84.86 84.03 S3R [36] 86.38 68.45 92.19 93.55 86.91 93.55 81.69 85.03 82.07 85.32 91.64 94.59 83.82 86.55 85.99 SSRL [13] 95.33 79.26 93.27 91.74 89.06 92.25 87.36 80.24 87.75 84.5 92.31 94.22 88.17 88.88 87.43 ActionCLIP [34] 91.88 90.47 89.21 86.87 81.31 94.08 83.23 94.34 82.82 70.53 91.60 94.06 89.89 87.71 82.30 AnomalyCLIP [42] 75.03 94.56 96.66 94.80 90.08 94.79 88.76 93.30 86.85 87.45 89.47 97.00 89.78 90.66 86.36 AnomalyCLIP [42] w/ Frame Restriction 69.51 (-5.52) 90.03 (-4.53) 92.44 (-4.22) 88.47 (-6.33) 82.38 (-7.7) 90.44 (-4.35) 85.83 (-2.93) 90.64 (-2.66) 82.22 (-4.63) 83.21 (-4.24) 87.43 (-2.04) 93.95 (-3.05) 83.48 (-6.3) 86.15 (-4.5) 83.35 (-3.01) MissionGNN (Ours) 93.68 92.95 94.15 96.89 92.31 94.79 92.38 97.16 81.40 90.57 92.94 97.74 84.03 92.38 84.48
TABLE III: Results of the state-of-the-art methods and MissionGNN on XD-Violence
Method Class mAP AP Abuse CarAccident Explosion Fighting Riot Shooting CLIP zero-shot [22] 0.32 12.21 22.26 25.25 66.60 1.26 21.32 27.21 RTFM [29] 9.25 25.36 53.53 61.73 90.38 18.01 43.04 77.81 S3R [36] 2.63 23.82 45.29 49.88 90.41 4.34 36.06 80.26 ActionCLIP [34] 2.73 25.15 55.28 58.09 87.31 12.87 40.24 61.01 AnomalyCLIP [42] 6.10 31.31 68.75 71.44 92.74 26.13 49.41 78.51 AnomalyCLIP [42] w/ Frame Restriction 4.2 (-1.9) 21.24 (-10.07) 48.34 (-20.41) 47.08 (-24.36) 73.33 (-19.41) 16.53 (-9.6) 35.12 (-14.29) 65.51 (-13) MissionGNN (Ours) 10.98 8.91 42.49 34.06 81.64 15.03 32.18 98.42
V Discussion
Significant Performance Drop of the Previous State-of-the-Art Model on XD-Violence in Real-Time Scenarios. The notable performance drop of AnomalyCLIP on the XD-Violence dataset in real-time scenarios stems from the dataset’s highly diverse and dynamic nature. XD-Violence features a wide range of video types, including fast-paced scenes from movies and sports events, where frequent changes in camera angles and scene transitions are common. These scene swaps make it unrealistic for real-time applications, as relevant frames may be separated by long periods within the video sequence. AnomalyCLIP performs well in this dataset because it processes entire video sequences, allowing it to consider both past and future frames, effectively capturing long-term relationships and subtle anomalies spread across disparate time segments. This capability is particularly advantageous in offline settings, where the model can access all frames at once, contributing to its superior performance in VAR tasks. However, in real-time applications, where future frames are not available, this reliance on long-term temporal context becomes a major limitation. When AnomalyCLIP is constrained to operate within a 30-frame window, as required for real-time analysis, it experiences a significant performance drop, since it can no longer rely on distant future or past frames to refine its predictions. In contrast, MissionGNN is designed specifically for real-time scenarios, where timely and efficient anomaly detection is crucial. By focusing on short-term relationships within the immediate frame window, MissionGNN is more suitable for real-world applications like violence alerting, where multiple scenes and unpredictable events make it impractical to rely on future information. Although MissionGNN may not capture long-term dependencies as effectively, its real-time responsiveness makes it more practical in scenarios requiring immediate detection. This comparison highlights the trade-off between models like AnomalyCLIP, optimized for comprehensive analysis with long-term context, and those like MissionGNN , which are tailored for efficient, real-time anomaly detection despite scene-swapping challenges in datasets like XD-Violence.
Limitations on mission-specific graph generation. Even though we introduce our GNN reasoning approach to deal with the heavy gradient computation from large pre-trained models, the efficacy of our framework heavily relies on the configuration of each mission-specific KG which can limit the performance of the model. Specifically, the construction of mission-specific KGs is facilitated by the output of GPT-4, which denotes the relevant terms associated with each mission. As a result, particularly in the context of VAR tasks, the selection of inappropriate terms by GPT-4 can significantly impair the model’s performance even with our reasoning GNN model. In contrast, AnomalyCLIP adopts gradient-based prompt optimization, which is advantageous for more effective related word vector extraction at the cost of necessitating extensive gradient computations. Thus, there exists a trade-off between the computational cost of prompt optimization. Future efforts will focus on mitigating the trade-off relationship by addressing the constraints of our mission-specific graph generation framework. This study will aid in tackling the trend where many LLMs are accessible solely through APIs, hindering direct access to their model weights.
Limitations on anomaly localization. Given that our decaying threshold-based anomaly localization method shown in Equation 7. alters the threshold statically, without taking into account the distribution of video frames, it becomes imperative to meticulously select the hyperparameter $\alpha_{d}^{iter}$ to achieve satisfactory test outcomes. In future works, it is essential to devise a mechanism that eliminates the necessity of employing this hyperparameter and instead enables dynamic adjustment of the threshold during model training.
VI Conclusions
In conclusion, our study introduces MissionGNN a novel hierarchical multimodal GNN-based approach for weakly-supervised video anomaly recognition, leveraging an automatically generated mission-specific knowledge graph to efficiently address the challenges of anomaly detection in video data. By circumventing the computational burdens from gradient computation of large multimodal models and eschewing fixed video segmentation, our model enables fully frame-level training and inference, demonstrating state-of-the-art performance on benchmark datasets while having training efficiency in terms of less GPU memory usage. This work not only marks a significant advancement in VAR but also offers a scalable, practical framework applicable to a broad range of real-world scenarios, paving the way for future innovations in intelligent surveillance and beyond.
Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, and Army Research Office Grant #W911NF2410360. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.